4 consecutive whole numbers such that the sum of the first three equals the fourth number
Answer by stanbon(50474)

4 consecutive whole numbers such that the sum of the first three equals the fourth number
----
1st: x-2
2nd:x-1
3rd:x
4th:x+1
----
sum = 3(x+1)
4x-2 = 3x+3
x = 5
---
1st: x-2 = 3
2nd: x-1 = 4
3rd: x = 5
4th: x+1 = 6
-----------------
cheers,
Stan H.